The radius of cross-section of a solid cylindrical rod of iron is 50cm the cylinder is melted down and formed into 6 solid spherical balls of the same radius as that of the cylinder the length of the rod(in metres)is
Answer by mananth(16946)

Radius of cylinder = 50 cm
length of cylinder = x
Volume of cylinder = pi r^2*x
radius of sphere = x
volume of sphere = 4/3 pi r^3
6 spheres
total volume of 6 spheres = 6*(4/3) pi r^3
8 pir^3
Volumes are equal
8 pi r^3= pi r^2 x
8r = x
r= 50
x= 8*50
400 cm length of rod